# Flow-Modulated Scoring for Semantic-Aware Knowledge Graph Completion

### v2

@@ 4.5. Efficiency Analysis: RQ3
We performed parametric volume analysis on both datasets, on the smaller dataset DDB14 as in Table 6 and on the larger dataset FB15k-237 as in Table 7. The model complexity is $\mathcal{O}(n\cdot K^{Hops})$. The result demonstrates that FMS is much more storage-efficient than embedding-based methods, since it does not need to calculate and store entity embeddings. Furthermore, FMS is significantly more effective than GNN-based methods such as PathCon. FMS can achieve superior performance with fewer hops and a smaller number of sampled neighbors. Additionally, it does not require explicitly modeling paths from the head entity to the tail entity like PathCon; instead, it employs flow-matching to learn a dynamic transition vector field, which represents semantically context-guided paths.
4.6. Case Study: RQ4
To quantitatively evaluate the impact of the Flow-Modulated Scoring (FMS) framework, we investigate using an example from the FB15k-237 dataset. The task is defined as: (California College of the Arts, ?, United States Dollar), for which the correct relations are: Revenue Currency, Operating Income Currency, Local Tuition Currency, Assets Currency, Domestic Tuition Currency, and Endowment Currency.  Figure 5 presents T-SNE (Van der Maaten and Hinton, 2008) visualizations to compare the representations learned by FMS against those from traditional static scoring methods.
Figure 5. T-SNE visualization of Static Score versus Flow-Modulated Score representations on FB15k-237.
@@ 4.6. Case Study: RQ4
Central to the figure, the ”Static Score Embedding” (black square) represents the generalized initial understanding that a traditional static model might offer for the association between “California College of the Arts” and “United States Dollar.” Due to its singular nature, this single point struggles to simultaneously and precisely encapsulate the specific role of “United States Dollar” as a currency across six different financial contexts. The power of FMS is evident in its subsequent dynamic processing. Observing the ’Dynamic Scores’ (colored stars) in the figure, we see that FMS is not content with this single, potentially ambiguous static representation. Instead, it actively transforms or “flows” from this static starting point to generate specific embedding representations for each potential relation. Each arrow in the figure symbolizes this context-driven modulation process, guiding the initial representation towards more relevant and fine-grained semantic space regions associated with specific relation types.
@@
FMS successfully distinguishes between multiple concurrently valid relations. It does not merely identify the generic concept of “currency”; rather, it can precisely differentiate specific relations such as “Revenue Currency” and “Assets Currency” within the embedding space. This clearly demonstrates its exceptional performance in capturing the dynamic nature of knowledge graph relations.
5. Conclusion
@@
In this work, we introduced Flow-Modulated Scoring (FMS), a novel framework designed to address the limitations of existing knowledge graph completion methods in modeling multifaceted semantic relationships and their dynamic nature. FMS innovatively combines a semantic context learning module, which generates context-aware entity embeddings, with a condition flow-matching module that learns the transformation dynamics between entities based on this context. By using the predicted vector field to dynamically modulate a base static score, FMS effectively synergizes rich static representations with conditioned dynamic flow information. Our comprehensive experiments on several well-established benchmarks demonstrate that FMS achieves state-of-the-art performance, validating its capability to capture more comprehensive relational semantics and advance the field of relation prediction.
6. GenAI Usage Disclosure
We confirm that no generative AI software was used to create any of the above content, including charts, code, and data. Any such AI tools were used solely to refine spelling, grammar, and translation to improve the readability of the paper.

### v3

@@ 4.5 Efficiency Analysis: RQ3
We performed parametric volume analysis on both datasets, on the smaller dataset DDB14 as in Table VII and on the larger dataset FB15k-237 as in Table VIII. The model complexity is $\mathcal{O}(n\cdot K^{Hops})$. The result demonstrates that FMS is much more storage-efficient than embedding-based methods, since it does not need to calculate entity embeddings. Furthermore, FMS is significantly more effective than GNN-based methods such as PathCon. FMS can achieve superior performance with fewer hops and a smaller number of sampled neighbors. Additionally, it does not require explicitly modeling paths from the head entity to the tail entity like PathCon; instead, it employs flow-matching to learn a dynamic evolution vector field, which represents semantically context-guided paths.
4.6 Case Study on Contextual Information: RQ4
To demonstrate the robust capability of our FMS in capturing the semantic context of relations, we present heatmaps of the modulated context-aggregated head/tail entity matrix on DDB14 (14 relation types) and WN18RR (11 relation types) in Figure 5. The intensity of the entries in Figure 5 represents the strength of the correlation between the presence of a contextual relation and a predicted relation. Relation IDs and their corresponding meanings for DDB14 and WN18RR are detailed in Table IX.
TABLE IX: Relation IDs and Meanings on DDB14 and WN18RR Datasets.
DDB14 Dataset
ID Relation Meaning ID Relation Meaning
0 belong(s) to the category of 7 interacts with
1 is a category subset of 8 belongs to the drug family of
2 may cause 9 belongs to drug super-family
3 is a subtype of 10 is a vector for
4 is a risk factor for 11 may be allelic with
5 is associated with 12 see also
6 may contraindicate 13 is an ingredient of
WN18RR Dataset
ID Relation Meaning ID Relation Meaning
0 hypernym 6 has_part
1 derivationally_related_form 7 member_of_domain_usage
2 instance_hypernym 8 member_of_domain_region
3 also_see 9 verb_group
4 member_meronym 10 similar_to
5 synset_domain_topic_of
Figure 5 reveals a sophisticated learning mechanism that moves beyond superficial pattern matching. The prevalence of strong off-diagonal correlations indicates that the model transcends simple pattern memorization—typified by rules like $(h,R,t^{\prime})\Rightarrow(h,R,t)$  [19]. Instead, the model constructs a network of latent semantic rules, which we categorize into the following core principles.
  • Domain-Specific Rule Induction: FMS demonstrates a remarkable ability to induce logical rules that are specifically adapted to the intrinsic properties of a given knowledge domain.
    In DDB14, it captures a clinical cause-and-effect relationship, inferring a rule that connects risk factors to clinical contraindications:
    $\small(h,\text{is\_risk\_factor\_for},t^{\prime})\Rightarrow(h,\text{may\_contraindicate},t).$
    In WN18RR, it models a fundamental linguistic principle linking etymology to semantics:
    $\small(h,\text{etymologically\_related},t^{\prime})\Rightarrow(h,\text{usage\_domain},t).$
  • Modeling of Hierarchical Structures: FMS effectively captures hierarchical knowledge structures, enabling it to perform complex, multi-level reasoning from specific instances to general categories.
    In DDB14, it masters compositional reasoning, inferring macroscopic class membership from microscopic constituents:
    $\small(h,\text{is\_ingredient\_of},t^{\prime})\Rightarrow(h,\text{belongs\_to\_drug\_family},t).$
    In WN18RR, it demonstrates a form of inductive inference, leveraging horizontal similarity to determine vertical class membership:
    $\small\begin{split}(h,\text{similar\_to},h^{\prime})\wedge(h^{\prime},\text{instance\_hypernym},t)\\ \Rightarrow(h,\text{instance\_hypernym},t).\end{split}$
4.7 Case Study on Dynamic Evolution: RQ5
To quantitatively evaluate the impact of the Flow-Modulated Scoring (FMS) framework, we investigate a case study from the FB15k-237 dataset. The task, detailed in Table X, involves predicting the multiple valid financial relationships between an arts college and a currency.
Figure 6 presents T-SNE [83] visualizations to compare the representations learned by FMS against those from traditional static scoring methods.
@@ 4.7 Case Study on Dynamic Evolution: RQ5
Central to the figure, the ”Static Score Embedding” (black square) represents the generalized initial understanding that a traditional static model might offer for the relation between ”California College of the Arts” and ”United States Dollar”. Due to its singular nature, this single point struggles to simultaneously and precisely encapsulate the specific role of ”United States Dollar” as a currency across six different financial contexts. The power of FMS is evident in its subsequent dynamic processing. Observing the ”Dynamic Scores” (colored stars) in the figure, we see that FMS is not content with this single, potentially ambiguous static representation. Instead, it actively flows from this static starting point to generate specific embedding representations for each potential relation. Each arrow in the figure symbolizes this context-driven modulation process, guiding the initial representation towards more relevant and fine-grained semantic space regions associated with specific relation types.
@@
FMS successfully distinguishes between multiple concurrently valid relations. It does not merely identify the generic concept of ”currency”; rather, it can precisely differentiate specific relations such as ”Revenue Currency” and ”Assets Currency” within the embedding space. This clearly demonstrates its exceptional performance in capturing the dynamic nature of knowledge graph relations.
TABLE X: Case study example from the FB15k-237 dataset.
Component Value
Head Entity California College of the Arts
Tail Entity United States Dollar
Query (California College of the Arts, ?, United States Dollar)
Ground-Truth Revenue Currency Operating Income Currency Local Tuition Currency Assets Currency Domestic Tuition Currency Endowment Currency
Figure 6: T-SNE visualization of Static Score versus Flow-Modulated Score representations on FB15k-237.
5 Entity Prediction Experiment
Building on its demonstrated superiority in relation prediction, we further validate FMS’s versatility and power on the fundamental task of entity prediction. As established previously, these two tasks are inter-reducible. To formally adapt our model for this task, we model the probability distribution over all candidate entities. Specifically, for a tail prediction query $(h,r,?)$, the probability of an entity $t$ is given by applying a SoftMax function over the scores derived from our model FMS:
$p(t|h,r)=\text{SoftMax}\left(s_{(h,t)}\right).$ (24)
Head prediction $(?,r,t)$ is handled symmetrically.
It is crucial to note that this adaptation necessitates a change in the training paradigm. Unlike in relation prediction where the target relation $r$ is masked from the input pair $(h,t)$, the model for entity prediction must be trained with the relation $r$ as part of the input condition to properly predict the target entity. This rigorous approach allows us to confirm whether the superior performance of FMS extends to this fundamental KG completion task.
TABLE XI: Transductive results. Performance comparison on FB15k-237, NELL995, Kinship and UMLS datasets. Best results are in bold, second best are underlined.
Type Model FB15k-237 NELL995 Kinship UMLS
MRR H@1 H@3 MRR H@1 H@3 MRR H@1 H@3 MRR H@1 H@3
Emb-based TransE 28.9 19.8 32.4 32.6 22.5 39.9 26.3 0.5 41.8 24.7 7.3 29.7
ComplEx 24.7 15.8 27.5 23.7 19.7 25.2 41.8 24.2 49.9 41.1 27.3 46.8
RotaE 33.7 24.1 37.5 35.6 29.8 39.5 65.1 50.4 75.5 74.4 63.6 82.2
ConvE 32.5 23.7 35.6 51.1 44.6 - 68.5 55.2 78.6 75.6 69.7 80.7
TDN 35.0 26.3 39.5 - - - 78.0 67.7 86.7 86.0 82.4 87.3
Rule-based RNNLogic 34.4 25.2 38.0 41.6 36.3 - 72.2 59.8 81.4 84.2 77.2 89.1
DRUM 23.8 17.4 26.1 53.2 46.0 - 33.4 18.3 37.8 54.8 35.8 69.9
GNN-based PathCon 50.6 39.6 56.3 49.5 44.4 51.9 78.8 65.6 91.5 90.1 85.7 93.2
NBFNet 41.5 32.1 45.4 52.5 45.1 - 60.6 43.5 72.5 77.8 68.8 84.0
RED-GNN 38.2 31.7 42.9 54.9 46.3 60.9 78.5 69.0 85.4 96.3 94.5 97.6
AdaProp 42.5 31.0 47.0 55.2 49.7 58.6 83.6 74.9 91.1 96.4 94.6 97.9
Diffusion-based FDM 48.5 38.6 52.9 - - - 83.7 76.1 89.7 92.2 89.3 94.4
KGDM 52.0 42.3 56.6 - - - 78.9 68.7 87.0 90.9 87.2 93.7
DiffusionE 37.6 29.4 - 52.2 49.0 - - - - 97.0 95.7 -
Ours FMS 67.2 57.4 72.9 60.6 54.9 65.8 99.3 98.7 99.9 92.5 89.0 94.8
5.1 Transductive Settings
5.1.1 Datasets
We conduct experiments on four knowledge graph datasets: (i) FB15k-237 (see Section 4.1 for details); (ii) NELL995, specifically an extended version (see Section 4.1 for background); (iii) Kinship, a knowledge graph describing kinship relations, focusing on family and blood relations [84]; (iv) UMLS, a biomedical knowledge graph encompassing medical terminology, concepts, and their interrelationships [85]. The statistics above are summarized in Table XIII.
5.1.2 Baselines
We categorize all baseline methods into four groups: (i) Embedding-based: TransE [20], ComplEx [40], RotatE [14], ConvE [71], TDN [86], FDM [37]; (ii) Rule-based: RNNlogic [87], DRUM [75]; (iii) GNN-based:Pathcon [19], NBFNet [88], RED-GNN [26], AdaProp [89]; (iv) Diffusion-Based: FDM [37], KGDM [47], DiffusionE [68].
5.1.3 Evaluation Metrics
To ensure a fair comparison, our training and evaluation setup follows that of RED-GNN [26]555 https://github.com/LARS-research/RED-GNN. For evaluation, each test triplet $(h,r,t)$ is used to create two entity prediction queries: a tail prediction query $(h,r,?)$ and a head prediction query $(?,r,t)$. The objective is to rank the ground-truth entity ($t$ for the former, $h$ for the latter) against a set of candidate entities. These candidates are generated by corrupting the triplet; for instance, in tail prediction, the ground truth $t$ is replaced by every other entity $t^{\prime}$ in the knowledge graph to form negative samples $(h,r,t^{\prime})$. We report Mean Reciprocal Rank (MRR) and Hits@N under the standard filtered setting [14], which ensures that any other valid triplets appearing in the candidate set are removed before ranking.
5.1.4 Implementation Details
Our experimental setup is largely consistent with the details provided in Section 4.1. However, we make specific adjustments to three key hyperparameters: the number of context hops, the number of neighbor samples, and the Top-K value. This modification is motivated by the significant disparity between the number of entities and relations; for instance, FB15k-237 contains 14,541 entities but only 237 relations. Modeling the probability distribution over a much larger set of entities necessitates richer semantic information. Therefore, we uniformly set the number of context hops to 2 and the Top-K value to 10 across all datasets. The number of neighbor samples is generally set to 16, with a specific exception for the NELL-995 dataset, for which we use a value of 10.
5.2 Inductive Settings
5.2.1 Datasets
Building upon this, we proceed with four subsets each from the FB15k237 and NELL-995 datasets, for a detailed examination of the dataset specifics, refer to Section 4.2.
5.2.2 Baselines
For inductive learning, traditional Embedding-based methods are not suitable for the task due to their inability to generalize to unseen entities, we choose the following two classes of classical and effective models: (i) Rule-based: RuleN [81], NeuralLP [82], DRUM [75]; (ii) GNN-based: GraIL [17], PathCon [19].
5.2.3 Baselines
Since completion methods that learn entity embeddings during training cannot work in this setting, we select baselines from the following three categories: (i) Rule-based methods: NeuralLP [82], DRUM [75], and RuleN [81]. (ii) GNN-based methods: GraIL [17], PathCon [19], NBFNet [88], RED-GNN [26], and AdaProp [89]. (iii) Diffusion-based methods: DiffusionE [68]. It is worth noting that KGDM and FDM use an embedding-based approach to encode entities and relations; therefore, they are not considered in the inductive learning setting.
5.2.4 Evaluation Metrics
The setup and parameter selection are consistent with the transduction entity prediction experiments.
TABLE XII: MRR comparison for entity prediction by multiple relation category on FB15k-237.
TransE RotatE NBFNet FDM FMS
Head Pred
1-N 7.9 8.1 16.5 20.3 39.8
N-1 45.5 46.7 49.9 55.9 69.0
N-N 22.4 23.4 34.8 42.3 67.7
Tail Pred
1-N 74.4 74.7 79.0 82.6 85.1
N-1 7.1 7.0 12.2 16.7 33.9
N-N 33.0 33.8 45.6 54.3 66.5
TABLE XIII: Statistics of all datasets. $\mathbb{E}[d]$ and $\mathrm{Var}[d]$ are mean and variance of the node degree distribution, respectively.
Dataset Nodes Relations Train Val Test $\mathbb{E}[d]$ $\mathrm{Var}[d]$
NELL995 74.5k 200 37.4k 543 2.8k 2.6 124.7
Kinship 104 25 3.2k 2.1k 5.3k 205.5 2.8
UMLS 135 46 1.3k 569 633 37.4 1205.3
TABLE XIV: Inductive results. The best results are shown in bold while the second-best results are shown underlined.
Type Model FB15k-237 NELL-995
v1 v2 v3 v4 v1 v2 v3 v4
MRR H@10 MRR H@10 MRR H@10 MRR H@10 MRR H@10 MRR H@10 MRR H@10 MRR H@10
Rule-based Neural-LP 32.5 46.8 38.9 58.6 40.0 57.1 39.6 59.3 61.0 87.1 36.1 56.4 36.7 57.6 26.1 53.9
DRUM 33.3 47.4 39.5 59.5 40.2 57.1 41.0 59.3 62.8 87.3 36.5 54.0 37.5 57.7 27.3 53.1
RuleN 36.3 44.6 43.3 59.9 43.9 60.0 42.9 60.5 61.5 76.0 38.5 51.4 38.1 53.1 33.3 48.4
GNN-based GraIL 27.9 42.9 27.6 42.4 25.1 42.4 22.7 38.9 48.1 56.5 29.7 49.6 32.2 51.8 26.2 50.6
PathCon 31.7 53.9 33.3 54.7 36.1 54.7 40.1 61.3 - - 37.3 59.4 45.6 59.0 29.2 44.8
NBFNet 30.7 51.7 36.9 63.9 33.1 58.8 30.5 55.9 58.4 79.5 41.0 63.5 42.5 60.6 28.7 59.1
RED-GNN 36.9 48.3 46.9 62.9 44.5 60.3 44.2 62.1 63.7 86.6 41.9 60.1 43.6 59.4 36.3 55.6
AdaProp 31.0 55.1 47.1 65.9 47.1 63.7 45.4 63.8 64.4 88.6 45.2 65.2 43.5 61.8 36.6 60.7
Diffusion-based DiffusionE 31.0 52.5 47.4 66.1 45.6 62.7 44.6 62.5 67.3 87.3 42.4 66.0 45.8 62.0 30.8 52.3
Ours FMS 42.3 63.3 52.8 70.6 53.5 73.3 57.6 75.6 - - 50.4 68.8 54.7 67.6 39.1 53.9
5.3 Overall Performance
To evaluate FMS on entity prediction, we conducted a parallel set of experiments, with results presented for both transductive and inductive settings.
The transductive results are summarized in Table XI. FMS demonstrates superior performance across the majority of datasets and metrics. Specifically, FMS achieves a remarkable 25.2% relative increase in MRR over the strongest baseline on FB15k-237, a 5.4% relative increase on NELL995, and an 15.6% relative increase on Kinship. Particularly on the Kinship dataset, FMS achieves a near-perfect MRR of 99.3, showcasing its exceptional modeling capacity on dense, complex graph structures. While DiffusionE shows the best performance on UMLS, FMS still delivers highly competitive results, establishing itself as a new state-of-the-art model for transductive entity prediction on diverse benchmarks.
To delve deeper into its capabilities, we analyze FMS’s performance on different relation categories within FB15k-237, as detailed in Table XII. FMS achieves overwhelmingly superior performance across all prediction scenarios (1-N, N-1, N-N for both Head and Tail prediction). The improvements are especially pronounced in more challenging categories. For instance, in the N-1 tail prediction task, FMS more than doubles the MRR of the best baseline, increasing it from 16.7 to 33.9. This uniform dominance indicates that FMS successfully captures the intricate and varied semantic mappings between entities, demonstrating its exceptional ability to handle complex reasoning scenarios across all relation types.
The inductive entity prediction results are presented in Table XIV. FMS consistently and significantly outperforms all baselines across all versions of the FB15k-237 and NELL-995 datasets, often by a large margin. For example, on FB15k-237 v4, FMS achieves an MRR of 57.6, surpassing the next-best model’s 45.4. Similarly, on NELL-995 v3, FMS’s MRR of 54.7 is substantially higher than the runner-up’s 45.8. Furthermore, the model’s strong and stable performance across different subgraph versions (v1 to v4) highlights its excellent robustness and generalization capabilities in unseen entity settings. This consistent and significant outperformance across diverse settings underscores the powerful generalization capabilities of FMS in inductive learning.
6 Conclusion
@@
In this work, we introduced Flow-Modulated Scoring (FMS), a novel framework designed to address the limitations of existing knowledge graph completion methods in modeling the semantic contextual dependency and dynamic nature of relations. FMS innovatively combines a semantic context learning module, which generates context-aware entity embeddings, with a condition flow-matching module that learns the evolution dynamics between entities based on this context. By using the predicted vector field to dynamically modulate a base static score, FMS effectively synergizes rich static representations with conditioned dynamic flow information. Our comprehensive experiments on several well-established benchmarks demonstrate that FMS achieves state-of-the-art performance, validating its capability to capture more comprehensive relational semantics and advance the field of knowledge graph completion.
Despite its empirical success, a primary limitation of FMS lies in its significant reliance on the semantic quality of relational features. The model’s performance is intrinsically tied to its ability to capture expressive relational patterns. Consequently, on large-scale knowledge graphs characterized by a vast and diverse set of relations—particularly where some relations may be sparse or semantically ambiguous—the efficacy of FMS may be constrained.
To address this limitation, a promising avenue for future research is the incorporation of entity-level features. By jointly modeling both entity and relational semantics, we can potentially achieve a more holistic and robust representation of factual knowledge. However, this extension introduces a non-trivial trade-off: integrating rich entity features would significantly increase the model’s parameter count and computational overhead. Therefore, a key challenge will be to strike a balance between enhanced expressiveness and model efficiency. This necessitates the exploration of more scalable encoding and aggregation mechanisms, such as parameter-sharing techniques or lightweight attention modules, to create a more comprehensive yet practical model.